1. Cyclopropane, C3H6, is an anesthetic given by inhalation.

A. How many moles of cyclopropane are there in 57 g of compound?

B. How many grams are in a 0.48 mole sample of cyclopropane?

C. In your answer in letter A, how many molecules are there in the compound?



2. Mepiridine hydrochloride, a sedative has a chemical formula of C15H21O2NHCl.

A. What is the molecular mass or molar mass of Mepiridine hydrochloride?

B. Compute the number of moles present in 106 grams of Mepiridine hydrochloride?

C. In your answer in letter B, what is the number of molecules present in the compound?



3. Platinol or Cisplatin has a chemical formula of PtCl2H6N2. It is a powerful drug used in the treatment of certain cancers.

A. Compute the molecular or molar mass of Platinol.

B. Calculate the number of moles in 285.3 grams of Platinol.

C. Compute the number of hydrogen atoms in 0.98 mol of platinol.
1
Expert's answer
2020-05-07T14:15:40-0400

1A: the number of moles can be found by dividing the mass of the substance by its molar weight.


"n=\frac{m}{\Mu}=\frac{57}{12\cdot3+6}=1.36\text{ mol.}"

1B:


"m=n\Mu=0.48\cdot(12\cdot3+6)=20.2\text{ g}."

1C: 1 mole of every substance contains 6.022x1023 molecules. Therefore, n moles contain


"N=N_A\cdot n=6.022\cdot10^{23}\cdot1.36=8.19\cdot10^{23}."

2A: The molecular mass, or molar mass (weight) is the sum of molar weights of all substances. All you need is to know the molar mass of each substance, multiply it by the number of the atoms of such kind, and add them:


"\Mu=12\cdot15+1\cdot21+16\cdot2+14+1+35.5=\\=283.5\text{ g/mol}."

2B:


"n=\frac{m}{\Mu}=\frac{106}{283.5}=0.374\text{ mol}."

2C:


"N=N_A\cdot n=6.022\cdot10^{23}\cdot0.374=2.25\cdot10^{23}."

3A:


"\Mu=195+35.5\cdot2+1\cdot6+14\cdot2=300\text{ g/mol}."

3B:


"n=\frac{m}{\Mu}=\frac{285.3}{300}=0.951\text{ mol}."

3C: first, find the number of Platinol molecules:

"N_P=N_{AP}\cdot n_P."

Next, from the chemical formula, we see that every single molecule of Platinol contains six hydrogen atoms. Therefore, the number of hydrogen atoms in nP=0.98 mol of platinol will be


"N_H=6N_P=6nN_{AP}=6\cdot0.98\cdot6.022\cdot10^{23}=\\=3.541\cdot10^{24}."